OpenStudy (anonymous):

how do i solve 3x+y=8

OpenStudy (unklerhaukus):

are you trying to solve for x? or for y?

OpenStudy (anonymous):

I don't really know what to do 3x+y=8 5x-y=8

OpenStudy (unklerhaukus):

oh, its simultaneous equations,

OpenStudy (anonymous):

y=2-3x/4

OpenStudy (unklerhaukus):

3x+y = 8 5x-y = 8 add the equations together (3x+y) + [5x-y]= (8) + [8] and simplify, the y-terms will drop out

OpenStudy (anonymous):

take 3x rght hand side

OpenStudy (unklerhaukus):

(3x+y) + [5x-y] = 3x+y + 5x-y = 3x+5x+y -y =

OpenStudy (anonymous):

okay I did that what now and is are yhu guys on here to help everyday

OpenStudy (anonymous):

is any body on here still

OpenStudy (unklerhaukus):

= 3x+5x+y -y = (3+5)x+(1-1)y =

OpenStudy (anonymous):

am I done now unklerhaukus

OpenStudy (unklerhaukus):

what does it simplify to ?

OpenStudy (anonymous):

umm is it 8x+y

OpenStudy (unklerhaukus):

almost, but (1-1) is not 1

OpenStudy (anonymous):

okay how about 8x+(-9) is that Is that it

OpenStudy (unklerhaukus):

what is one, take away one?

OpenStudy (anonymous):

okay that is 0 so I get 8x+0 right

OpenStudy (unklerhaukus):

yeah (3x+y) + [5x-y] = 3x+y + 5x-y = 3x+5x+y -y = 8x+0

OpenStudy (unklerhaukus):

so you have 8x = 8+8 can you solve this for x?

OpenStudy (anonymous):

okay I divide by 8 right and then I get x=-8 am I right

OpenStudy (unklerhaukus):

8x = 8+8 if you divide by 8 you'll get x = 1 +1

OpenStudy (anonymous):

okay my answer is x=2 right

OpenStudy (unklerhaukus):

yes

OpenStudy (unklerhaukus):

good, now we have x, we can find y,

OpenStudy (unklerhaukus):

lets plug in x = 2 into that first equation 3x+y = 8 2(2) +y =8 solve this for y

OpenStudy (anonymous):

is y 2 thanks a lot

OpenStudy (unklerhaukus):

**whoops 3(2) +y =8

OpenStudy (anonymous):

okay does y = 2 then

OpenStudy (unklerhaukus):

yes!

OpenStudy (unklerhaukus):

Lets check our solution we got x=2, y=2 3x+y = 8 5x-y = 8 if we these in both these equations should be true 3(2)+(2) = 8 5(2)-(2) = 8 are these both right?

OpenStudy (anonymous):

yes they are are we done now

OpenStudy (unklerhaukus):

yep, good work

OpenStudy (anonymous):

okay THANK YHU A LOT I REALLY NEED THE EXTRA HELP IN MATH BYE

OpenStudy (unklerhaukus):

write your solution as either x=2, y=2 or (x, y) = (2, 2)

OpenStudy (unklerhaukus):

bye : )